# The description of two-dimensional altermagnetism

### v2

@@ III SPIN-MOMENTUM LOCKING IN MATERIAL CANDIDATE
For the spin layer group ${}^{1}4/^{1}m^{2}m^{2}m$, we have identified VP2H8(NO4)2, which has a layer group NO.56, as exhibiting this symmetry. The crystal structure of VP2H8(NO4)2 is schematically illustrated in FIG.3(a), with red and blue colors representing the real-space sublattices with opposite spins in altermagnetism. Being a van der Waals material, it can be exfoliated from the bulk compound. We obtain that the energy of the altermagnetic state is 0.47 meV per magnetic atom lower than that of the ferromagnetic state by DFT calculations indicating that its magnetic ground state is an altermagnetic state. The sublattices can be related by reflection through the xz-plane (yz-plane, (110) or ($01\overline{1}$ 0) plane), and a 4-fold rotation around z-axis maps sublattice to itself. Consequently, the spin-momentum locking of VP2H8(NO4)2 is determined by the symmetries $[C_{2}||\{m_{x},m_{y},m_{110},m_{1\overline{1}0}\}]$. That is the spin-degeneracy nodal line at $k_{x}=0$ and $k_{y}=0$ are protected by $[C_{2}||m_{x}]$ and $[C_{2}||m_{y}]$. The k-paths, related by $m_{x}$ ($m_{y},m_{110},m_{1\overline{1}0}$) but not separated by a reciprocal lattice vector, have opposite spin signs. FIG.3(b) and (c) show the DFT calculated band structure without SOC and the employed k-paths is shown in FIG.3(d). If we choose high-symmetry path along $\Gamma-X-M-\Gamma$ to calculate the band structure, only the spin-degeneracy bands are obtained, as illustrated in FIG.3(b), which is protected by symmetry $m_{x}$ and $m_{1\overline{1}0}$. However, the band structure of k-path along X1-$\Gamma$-X2 is spin-spliting and $\Gamma$-X1 has opposite spin to $\Gamma$-X2, which is a typical altermagnetic characteristic. Again, these results are consistent with the conclusions we have reached above.
@@ Acknowledgements.
This work is financially supported by National Natural Science Foundation of China (Grant No. 12074126). The computer times at the High Performance Computational center at South China University of Technology are gratefully acknowledged. YJZ is grateful to the valuable suggestions provided by Prof. Jiuwei Liu.

### v3

@@ III SPIN-MOMENTUM LOCKING IN MATERIAL CANDIDATE
For the spin layer group ${}^{1}4/^{1}m^{2}m^{2}m$, we have identified VP2H8(NO4)2 as exhibiting this symmetry. It has a layer group NO.56, which contains a 4-fold rotation symmetry around z-axis, a mirror symmetry through (110)plane and three glide reflections through xz-plane, yz-plane and ($01\overline{1}$ 0) plane. The crystal structure of VP2H8(NO4)2 is schematically illustrated in FIG.4(a), with red and blue colors representing the real-space sublattices with opposite spins in altermagnetism. Being a van der Waals material, it can be exfoliated from the bulk compound. We obtain that the energy of the altermagnetic state is 0.47 meV per magnetic atom lower than that of the ferromagnetic state by DFT calculations indicating that its magnetic ground state is an altermagnetic state. The sublattices can be related by reflection through the xz-plane (yz-plane, (110) or ($01\overline{1}$ 0) plane), and a 4-fold rotation around z-axis maps sublattice to itself. Consequently, the spin-momentum locking of VP2H8(NO4)2 is determined by the symmetries $[C_{2}||\{m_{x},m_{y},m_{110},m_{1\overline{1}0}\}]$. That is the spin-degeneracy nodal line at $k_{x}=0$ and $k_{y}=0$ are protected by $[C_{2}||m_{x}]$ and $[C_{2}||m_{y}]$. The k-paths, related by $m_{x}$ ($m_{y},m_{110},m_{1\overline{1}0}$) but not separated by a reciprocal lattice vector, have opposite spin signs. FIG.4(b) and (c) show the DFT calculated band structure without SOC and the employed k-paths is shown in FIG.4(d). If we choose high-symmetry path along $\Gamma-X-M-\Gamma$ to calculate the band structure, only the spin-degeneracy bands are obtained, as illustrated in FIG.4(b), which is protected by symmetry $m_{x}$ and $m_{1\overline{1}0}$. However, the band structure of k-path along X1-$\Gamma$-X2 is spin-spliting and $\Gamma$-X1 has opposite spin to $\Gamma$-X2, which is a typical altermagnetic characteristic. Again, these results are consistent with the conclusions we have reached above.
@@ Acknowledgements.
This work is financially supported by National Natural Science Foundation of China (Grant No. 12074126). The computer times at the High Performance Computational center at South China University of Technology are gratefully acknowledged. YJZ is grateful to the valuable suggestions provided by Prof. Junwei Liu.
@@ Appendix A COMPUTATIONAL DETAILS
Figure 5: The band structure with and without SOC of (a)RuF4, (b)Mn2P2S3Se3, (c)(f) MnTeMoO6, and (d)(e)VP2H8(NO4)2. The red and blue dotted lines represent the opposite channels, and the black solid lines represent band structure with SOC. (a) corresponds to FIG.2(b). (b) corresponds to FIG.2(e). (c) and (f) corresponds to FIG.3(c) and (d). (d) and (e) corresponds to FIG.4(b) and (c).
Appendix B COMPARISON OF BAND STRUCTURES WITH AND WITHOUT SOC
As is known to all, the absence of SOC effect is an approximation. In this appendix, we provide comparison of band structures with and without SOC for all materials, as illustrated in FIG.5. It is implied that the magnitude of spin-splitting originating from SOC effect is small than that arises from altermagnetism. Therefore, altermagnetism plays a key role in spin-splitting in these materials and the discussion above without SOC is appropriate. It is worth noting that the intersections of two bands in FIG.3(d) vanish when we consider SOC effect, as illustrated in FIG.5(f). Consequently, we conclude that these points are not Weyl points.